# Multi-modal Attribute Prompting for Vision-Language Models

### v2

@@ Abstract
Large pre-trained Vision-Language Models (VLMs), like CLIP, exhibit strong generalization ability to downstream tasks but struggle in few-shot scenarios. Existing prompting techniques primarily focus on global text and image representations, yet overlooking multi-modal attribute characteristics. This limitation hinders the model’s ability to perceive fine-grained visual details and restricts its generalization ability to a broader range of unseen classes. To address this issue, we propose a Multi-modal Attribute Prompting method (MAP) by jointly exploring textual attribute prompting, visual attribute prompting, and attribute-level alignment. The proposed MAP enjoys several merits. First, we introduce learnable visual attribute prompts enhanced by textual attribute semantics to adaptively capture visual attributes for images from unknown categories, boosting fine-grained visual perception capabilities for CLIP. Second, the proposed attribute-level alignment complements the global alignment to enhance the robustness of cross-modal alignment for open-vocabulary objects. To our knowledge, this is the first work to establish cross-modal attribute-level alignment for CLIP-based few-shot adaptation. Extensive experimental results on 11 datasets demonstrate that our method performs favorably against state-of-the-art approaches.
Keywords:
Vision-Language Model Few-Shot Adaptation
Copyright © 20xx IEEE. Personal use of this material is permitted. However, permission to use this material for any other purposes must be obtained from the IEEE by sending an email to pubs-permissions@ieee.org.
1 Introduction
@@
Large-scale pre-trained Vision-Language Models (VLMs), such as CLIP [38] and ALIGN [17], have demonstrated promising generalization power and transferability on a wide range of downstream tasks, including image classification [38], object detection [24, 2, 20] and 3D understanding [36, 51, 42]. Through contrastive training on a large-scale dataset of image-text pairs, CLIP achieves a global alignment between images and textual descriptions by learning a joint embedding space. The robust cross-modal alignment empowers the CLIP model with the open-vocabulary visual recognition capability. In CLIP, class-specific weights for open vocabulary classification can be constructed by plugging the class name in a predefined prompt template like ‘A photo of a [CLASS].’ Despite its impressive generalization capability, it remains challenging to adapt CLIP to downstream tasks in few-shot scenarios. The extensive size of the CLIP model and the limited training data available make it unfeasible to fine-tune the complete model for use in downstream few-shot tasks.
@@
The contributions of our method could be summarized in three-fold: (1) We propose Multi-modal Attribute Prompting, which jointly explores textual attribute prompting, visual attribute prompting, and attribute-level alignment between images and text categories. (2) We enhance fine-grained visual perception ability by modeling visual attribute features with visual attribute prompts. Moreover, we introduce attribute-level alignment, complementing global alignment, to achieve multi-level robust alignment between images and text categories. To our knowledge, this is the first work to model visual attributes and establish attribute-level alignment between images and text categories for adapting the pre-trained CLIP model to downstream few-shot tasks. (3) Extensive experimental results on 11 benchmark datasets demonstrate that our method performs favorably against state-of-the-art approaches.
2 Related work
@@
Pre-trained Vision-Language Models. In recent years, pre-trained vision-language models [38, 17, 41, 48, 47], have shown exceptional performance in diverse downstream tasks. Among them, CLIP stands out as a representative approach. By training its vision and text encoders to map both modalities closely in a shared embedding space, CLIP establishes a comprehensive global alignment between images and their corresponding textual descriptions, enabling open-vocabulary classification tasks. The classification result can be obtained by computing the similarity scores of the global image feature with class names encoded by the text encoder. However, as classification relies solely on the global matching score, the accuracy may be affected by disruptions in images, such as complex backgrounds, especially in few-shot settings. To improve the robustness of cross-modal alignment, we achieve multi-level alignment for CLIP by introducing additional attribute-level alignment between dynamically learned textual and visual attribute features. In this manner, our method enhances the fine-grained perception capability with the pre-trained global knowledge preserved.
@@
Prompt Learning. Prompt learning is initially introduced in the field of natural language processing (NLP) [37, 40, 19, 27, 26, 13, 28]. With language models frozen, prompt learning methods effectively facilitate the adaptation of pre-trained language models to downstream few-shot tasks by involving additional hand-crafted or learnable prompt tokens. Prompt learning has recently been employed to enhance the adaptation of the CLIP model to downstream few-shot tasks, where limited training samples are available. CoOp [5] constructs prompts by concatenating learnable continuous vectors and class name tokens. CoCoOp [50] extends CoOp by further learning a lightweight neural network to generate an input-conditional vector for each image, tackling the poor generalizability to broader unseen classes in CoOp [5]. ProDA [29] optimizes a set of prompts by learning the distribution of prompts. Instead of focusing on text-modal prompts, VPT [18] introduces learnable vectors to the Vision Transformer [10] to refine image features within the frozen vision encoder. DAPT [6], RPO [25], and MaPLe [21] improve the generalization ability of VLMs via multimodal prompting. These methods rely solely on class names for text prompt construction and may struggle to fully encapsulate categorical semantics.
@@
Textual Attribute Prompts. To enrich the semantic description for different classes, recent works [32, 31, 12], instead of relying solely on class names, have shifted towards the utilization of attribute descriptions to construct textual attribute prompts for each class. This shift is facilitated by the development of pre-trained large language models (LLMs) like the GPT family [4, 34]. Attribute descriptions can be easily obtained by querying the LLM with suitable question templates. However, these methods focus on attributes in text space only, neglecting the modeling of visual attributes, leading to limited visual perception capabilities of the model and misalignment between global visual and local textual features. In contrast, we jointly model visual and textual attribute features and establish attribute-level alignment between images and text categories.
3 Method
@@
In this section, we first provide a concise overview of CLIP [38]. Then, we present a comprehensive introduction to our proposed multi-modal attribute prompting, as illustrated in Figure 3, including textual attribute prompting, visual attribute prompting, and attribute-level alignment.
3.1 Review of CLIP
@@
3.2 Textual Attribute Prompting
@@
Textual Attribute Prompt Construction. We formulate $N$ textual attribute prompts for each class by combining attribute description sentences with a standardized prompt template. For instance, for the $k$-th class, with the template ‘A photo of a [CLASS]’ we construct a textual attribute prompt: $p_{k}^{n}$={A photo of a class ($k$), $t_{n,k}$}, where class ($k$) denotes the class name corresponding to the $k$-th class, and $t_{n,k}$ denotes the $n$-th attribute description for the $k$-th class. To enhance the adaptability of textual attribute prompts, we replace the hand-crafted context, i.e., ‘A photo of a’ with several learnable context vectors. By feeding the textual attribute prompts into the text encoder $\theta$, we can obtain encoded textual attribute prompts:
$\boldsymbol{G}_{k}=\{{g}_{k}^{n}|_{n=1}^{N}\},g_{k}^{n}=\theta(p_{n}^{k}),$ (2)
@@
3.3 Visual Attribute Prompting
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] indicates the concatenation along the sequence length dimension. In early layers of ViT, the visual attribute prompts progressively aggregate image regional features through interaction with image tokens.
@@
[…]
@@
3.4 Attribute-Level Alignment
To achieve precise alignment between visual attribute prompts $\{{u_{i}}^{(L)}\}_{i=1}^{M}$ and textual attribute prompts $\boldsymbol{G_{k}}=\{g_{k}^{n}|_{n=1}^{N}\}$, we formulate the attribute-level matching task as an Optimal Transport (OT) problem [43]. For simplicity, we refer to $\{{u_{i}}^{(L)}\}_{i=1}^{M}$ as $\boldsymbol{F}=\{f_{m}|_{m=1}^{M}\}$ hereafter. OT aims to find the optimal transportation plan with the minimal cost between two discrete distributions $\mu\in\mathbb{R}^{M}$, $\nu\in\mathbb{R}^{N}$. The optimal transportation plan $\mathbf{T}^{*}$ is obtained by minimizing the transportation cost:
$\begin{gathered}\mathbf{T}^{*}=\underset{\mathbf{T}\in\Pi(\mu,\nu)}{\arg\min}\langle\mathbf{T},\mathbf{C}\rangle,\\ \text{ s.t. }\quad\mathbf{T}\mathbf{1}=\mu,\mathbf{T}^{T}\mathbf{1}=\nu,\end{gathered}$ (7)
where $\prod(\mu,\nu)$ is the joint distribution with marginals $\mu$ and $\nu,\langle\cdot,\cdot\rangle$ denotes the cosine similarity, and $\mathbf{C}\in\mathbb{R}^{M\times N}$ represents the cost matrix of transporting $\mu$ to $\nu$. The Problem in Equation  (7) can be efficiently solved by the Sinkhorn algorithm [8]. To apply OT to the multi-modal attribute matching task, we formulate the attribute prompt sets as discrete uniform distributions. The cost matrix $\mathbf{C}$ is defined as the distances between attribute prompts. By solving Equation  (7), we can obtain $\mathbf{T}^{*}$ to serve as the alignment matrix, and then define the final similarity score between the visual attribute prompts $\boldsymbol{F}$ and textual attribute prompts $\boldsymbol{G}_{k}$ as:
$\psi(\boldsymbol{F},\boldsymbol{G}_{k})=\sum_{m=1}^{M}\sum_{n=1}^{N}<f_{m},g_{k}^{n}>\mathbf{T}_{mn}^{*},$ (8)
@@ 3.5 Training Objectives
where $B$ is the batch of image-text pairs, and $y_{i}$ denotes the ground-truth of $x_{i}$.
4 Experiments
We evaluate MAP in four settings: (1) generalization from base to novel classes within a dataset; (2) few-shot image classification; (3) domain generalization; (4) cross-dataset evaluation. All models used are based on the open-source CLIP [38].
Datasets. We use the 11 image recognition datasets following the setup in CoOp [5] for the first two settings. The benchmark includes Food101 (Foo) [3], DTD [7], Imagenet (Img) [9], Caltech101 (Cal) [11], EuroSAT (Eur) [14], StanfordCars (Car) [23], FGVCAircraft (FGV) [30], Flowers102 (Flo) [33], OxfordPets (Pet) [35], UCF101 (UCF) [35], and SUN397 (SUN) [46]. In the domain generalization setting, we utilize ImageNet as the source dataset, and its distinct domain variants, namely ImageNet-R (-R) [15], ImageNet-A (-A) [16], ImageNetV2 (V2) [39], and ImageNet-Sketch (-S) [44] as target datasets.

### v3

@@ Abstract
Pre-trained Vision-Language Models (VLMs), like CLIP, exhibit strong generalization ability to downstream tasks but struggle in few-shot scenarios. Existing prompting techniques primarily focus on global text and image representations, yet overlooking multi-modal attribute characteristics. This limitation hinders the model’s ability to perceive fine-grained visual details and restricts its generalization ability to a broader range of unseen classes. To address this issue, we propose a Multi-modal Attribute Prompting method (MAP) by jointly exploring textual attribute prompting, visual attribute prompting, and attribute-level alignment. The proposed MAP enjoys several merits. First, we introduce learnable visual attribute prompts enhanced by textual attribute semantics to adaptively capture visual attributes for images from unknown categories, boosting fine-grained visual perception capabilities for CLIP. Second, the proposed attribute-level alignment complements the global alignment to enhance the robustness of cross-modal alignment for open-vocabulary objects. To our knowledge, this is the first work to establish cross-modal attribute-level alignment for CLIP-based few-shot adaptation. Extensive experimental results on 11 datasets demonstrate that our method performs favorably against state-of-the-art approaches.
Index Terms:
Few-shot classification, Prompt learning, Vision-language model, Attribute.
I Introduction
Pre-trained Vision-Language Models (VLMs), such as CLIP [1] and ALIGN [2], have demonstrated promising generalization power and transferability on a wide range of downstream tasks [3, 4, 5, 6, 7, 8, 9], including image classification [1], object detection [10, 11] and 3D understanding [12, 13, 14]. Through contrastive training on a large-scale dataset of image-text pairs, CLIP achieves a global alignment between images and textual descriptions by learning a joint embedding space. The robust cross-modal alignment empowers the CLIP model with the open-vocabulary visual recognition capability. In CLIP, class-specific weights for open vocabulary classification can be constructed by plugging the class name in a predefined prompt template like ‘A photo of a [CLASS].’ Despite its impressive generalization capability, it remains challenging to adapt CLIP to downstream tasks in few-shot scenarios. Due to the large number of parameters in CLIP and the limited number of samples in few-shot task settings, naive fine-tuning of the entire model would likely lead to overfitting, resulting in performance degradation [15, 16].
@@ I Introduction
Our main contributions can be summarized as follows:
  • We propose Multi-modal Attribute Prompting, which jointly explores textual attribute prompting, visual attribute prompting, and attribute-level alignment between images and text categories. To our knowledge, this is the first work to model visual attributes and establish attribute-level alignment between images and text categories for adapting the pre-trained CLIP model to downstream few-shot tasks.
  • Extensive experimental results on 11 benchmark datasets demonstrate that our method performs favorably against state-of-the-art approaches.
II Related Works
@@
In recent years, pre-trained vision-language models [3, 4, 32, 33, 34, 35, 36] have shown exceptional performance in diverse downstream tasks. Among them, CLIP [1] stands out as a representative approach. By training its vision and text encoders to map both modalities closely in a shared embedding space, CLIP establishes a comprehensive global alignment between images and their corresponding textual descriptions, enabling open-vocabulary classification tasks. The classification result can be obtained by computing the similarity scores of the global image feature with class names encoded by the text encoder. However, as classification relies solely on the global matching score, the accuracy may be affected by disruptions in images, such as complex backgrounds, especially in few-shot settings [37, 38, 39, 40, 41, 42, 43], where only a few training samples are available. To improve the robustness of cross-modal alignment, we achieve multi-level alignment for CLIP by introducing additional attribute-level alignment between dynamically learned textual and visual attribute features. In this manner, our method enhances the fine-grained perception capability with the pre-trained global knowledge preserved.
II-B Prompt Learning.
@@
Prompt learning is initially introduced in the field of natural language processing (NLP) [44, 45, 46, 47, 48]. With language models frozen, prompt learning methods effectively facilitate the adaptation of pre-trained language models to downstream few-shot tasks by involving additional hand-crafted or learnable prompt tokens. Prompt learning has recently been employed to enhance the adaptation of the CLIP model to downstream few-shot tasks, where limited training samples are available. CoOp [16] constructs prompts by concatenating learnable continuous vectors and class name tokens. CoCoOp [18] extends CoOp by further learning a lightweight neural network to generate an input-conditional vector for each image, tackling the poor generalizability to broader unseen classes in CoOp [16]. ProDA [21] optimizes a set of prompts by learning the distribution of prompts. Instead of focusing on text-modal prompts, VPT [49] introduces learnable vectors to the Vision Transformer [50] to refine image features within the frozen vision encoder. DAPT [19], RPO [22], and MaPLe [23] improve the generalization ability of VLMs via multimodal prompting. PromptSRC [20] introduces regularization loss to prompt learning. These methods rely solely on class names for text prompt construction and may struggle to fully encapsulate categorical semantics.
II-C Textual Attribute Prompts.
@@
To enrich the semantic description for different classes, recent works [25, 26, 24], instead of relying solely on class names, have shifted towards the utilization of attribute descriptions to construct textual attribute prompts for each class. This shift is facilitated by the development of pre-trained large language models (LLMs) like the GPT family [27, 28]. Attribute descriptions can be easily obtained by querying the LLM with suitable question templates. However, these methods focus on attributes in text space only, neglecting the modeling of visual attributes, leading to limited visual perception capabilities of the model and misalignment between global visual and local textual features. In contrast, we jointly model visual and textual attribute features and establish attribute-level alignment between images and text categories.
II-D Visual Attributes.
Visual attributes refer to intuitive properties of objects, encompassing low-level semantics (e.g., color, texture, and shape) and high-level semantics (e.g., head, body, and tail of objects) [51]. Utilizing visual attributes has led to significant progress in various vision tasks, including image search [52], image recognition [53], and scene understanding [54]. Some previous works on learning attributes [55, 52, 56] usually require extensive manual attribute annotations, which are labor-intensive. Dealing with this issue, a recent work [57] developed an encoder-decoder network to unsupervisedly distill high-level attribute-specific vectors without requiring attribute annotations. VAPNet [58] achieves semantic details by utilizing local image patches to distill visual attributes from these discovered semantics. Different from these methods, our approach uniquely leverages visual prompts to model visual attributes. By incorporating visual attribute prompts as learnable tokens within Vision Transformers, our method captures and aggregates relevant image features effectively.
III Methodology
@@
In this section, we first provide a concise overview of CLIP [1]. Then, we present a comprehensive introduction to our proposed multi-modal attribute prompting, as illustrated in Figure 3, including textual attribute prompting, visual attribute prompting, and attribute-level alignment. The main symbols and instructions are shown in Table I.
TABLE I: Main symbols and instructions
Symbol Instruction
$\phi(\cdot)$ the image encoder
$\theta(\cdot)$ the text encoder
$\mathcal{V}$ set of class names
$\mathcal{C}$ the number of class names
$x$ the input image
$y$ the ground-truth label
$f$ the global image feature
$p_{k}^{n}$ the $n$-th textual attribute prompt of $k$-th class
$g_{k}^{n}$ encoded $n$-th textual attribute prompt of $k$-th class
$\boldsymbol{G}_{k}$ encoded textual attribute prompts of the $k$-th class
$l_{j}$ the $j$-th ViT layer
$E_{j}$ image tokens output from $j$-th ViT layer
$s_{j}$ [CLS] token output from $j$-th ViT layer
$U_{j}$ visual attribute prompts output from $j$-th ViT layer
$\boldsymbol{F}$ visual attribute prompts output from ViT
$T^{*}$ the optimal transportation plan
$\mathbf{\Gamma}$ adaptive visual attribute enhancement module
$\psi(\cdot,\cdot)$ similarity function
$M$ the number of visual attribute prompts
$N$ the number of textual attribute prompts
$L$ the number of transformer layers in ViT
$\boldsymbol{Q}$, $\boldsymbol{K}$, $\boldsymbol{V}$ queries, keys, and values in the attention layer
$W_{Q}$,$W_{K}$, $W_{V}$ linear projections of the attention layer
$\mathbf{1}_{N}$ N-dimensional all-one vector
$p$,$q$ discrete distributions
$\mu$,$\nu$ discrete probability vectors
III-A Review of CLIP
@@
III-B Textual Attribute Prompting
@@
Textual Attribute Prompt Construction. We formulate $N$ textual attribute prompts for each class by combining attribute description sentences with a standardized prompt template. For instance, for the $k$-th class, with the template ‘A photo of a [CLASS]’ we construct a textual attribute prompt: $p_{k}^{n}$={A photo of a class ($k$), $t_{k}^{n}$}, where class ($k$) denotes the class name corresponding to the $k$-th class, and $t_{k}^{n}$ denotes the $n$-th attribute description for the $k$-th class. To enhance the adaptability of textual attribute prompts, we replace the hand-crafted context, i.e., ‘A photo of a’ with several learnable context vectors. Following CoOp [16], we use four learnable class-agnostic context vectors, concatenated with the class name and attribute description to construct the textual attribute prompt. These vectors are optimized during training to better adapt to downstream tasks, providing a more flexible context.
By feeding the textual attribute prompts into the text encoder $\theta$, we can obtain encoded textual attribute prompts:
$\boldsymbol{G}_{k}=\{{g}_{k}^{n}|_{n=1}^{N}\},g_{k}^{n}=\theta(p_{k}^{n}),$ (2)
@@
III-C Visual Attribute Prompting
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] indicates the concatenation along the sequence length dimension. In early layers of ViT, the visual attribute prompts progressively aggregate image regional features through interaction with image tokens facilitated by the attention mechanism. Learnable visual attribute prompts compute similarity with image tokens and aggregate information accordingly. Similar to the [CLS] token in models like BERT [60] and ViT [50], visual prompts can read and aggregate visual information from image tokens [22]. Previous research [61, 62] indicates that ViTs will attend to local information in early layers. This property, together with the attention mechanism, helps aggregate image regional features.
@@
To better inject the semantic clues of selected textual prompts into visual attribute prompts, we design an attribute-aware cross-attention layer in $\mathbf{\Gamma}$. Here, the visual attribute prompt tokens $\{{u_{i}}^{(l)}\}_{i=1}^{M}$ function as queries $\boldsymbol{Q}$. Simultaneously, the textual attribute prompt features $\boldsymbol{G^{\prime}}$ of candidate classes are utilized as keys $\boldsymbol{K}$ and values $\boldsymbol{V}$. The enhanced visual attribute prompt $\tilde{u}_{i}^{(l)}$ is formulated as
@@
III-D Attribute-Level Alignment
To achieve precise alignment between visual attribute prompts $\{{u_{i}}^{(L)}\}_{i=1}^{M}$ and textual attribute prompts $\boldsymbol{G_{k}}=\{g_{k}^{n}|_{n=1}^{N}\}$, we formulate the attribute-level matching task as an Optimal Transport (OT) problem [30]. For simplicity, we refer to $\{{u_{i}}^{(L)}\}_{i=1}^{M}$ as $\boldsymbol{F}=\{f_{m}|_{m=1}^{M}\}$ hereafter. Optimal Transport (OT) [30] is a powerful tool to measure the distance between two distributions. Given two sets of feature points $\boldsymbol{F}=\{f_{m}|_{m=1}^{M}\}$ and $\boldsymbol{G_{k}}=\{g_{k}^{n}|_{n=1}^{N}\}$, their distributions can be formulated as $p=\sum_{m=1}^{M}\mu_{m}\delta_{f_{m}}$, $q=\sum_{n=1}^{N}\nu_{n}\delta_{g_{k}^{n}}$, $\delta_{f_{m}}$ is a Dirac delta function centered at a specific point $f_{m}$ in the embedding space. Here, $\mu\in\mathbb{R}^{M}$, $\nu\in\mathbb{R}^{N}$ are two discrete distribution vectors. We define the cost matrix between $\boldsymbol{F}=\{f_{m}|_{m=1}^{M}\}$ and $\boldsymbol{G_{k}}=\{g_{k}^{n}|_{n=1}^{N}\}$ as $\mathbf{C}\in\mathbb{R}^{M\times N}$, where $\mathbf{C}_{m,n}=1-\langle f_{m},g_{k}^{n}\rangle$ is the transport cost from $f_{m}$ to $g_{k}^{n}$. The transport cost between $p$ and $q$ is $\langle\mathbf{T},\mathbf{C}\rangle$, where $\mathbf{T}$ is the transport plan, and $\mathbf{T}_{m,n}$ is the probability or “flow” of transporting from $f_{m}$ to $g_{k}^{n}$. The goal of OT is to transport $p$ to $q$ at the smallest cost with the optimal transport plan $\mathbf{T}^{*}$:
$\begin{gathered}\mathbf{T}^{*}=\underset{\mathbf{T}\in\Pi(p,q)}{\arg\min}\langle\mathbf{T},\mathbf{C}\rangle,\\ \text{ s.t. }\quad\mathbf{T}\mathbf{1}_{N}=\mu,\mathbf{T}^{T}\mathbf{1}_{M}=\nu,\end{gathered}$ (7)
[…]
$\begin{gathered}\mathbf{T}^{*}=diag(U(t))\mathbf{A}diag(V(t)),\\ \mathbf{A}=exp(-\mathbf{C}/\gamma)\end{gathered}$ (8)
where in the $t$-th iteration, $U(t)=\mu/(\mathbf{A}V(t-1))$, $V(t)=\nu/\mathbf{A}^{T}U(t))$, with the initiation $V(0)=\mathbf{1}$. With Equation  (8), we can obtain $\mathbf{T}^{*}$ to serve as the alignment matrix, and then define the final similarity score between the visual attribute prompts $\boldsymbol{F}$ and textual attribute prompts $\boldsymbol{G}_{k}$ as:
$\psi(\boldsymbol{F},\boldsymbol{G}_{k})=\sum_{m=1}^{M}\sum_{n=1}^{N}\langle f_{m},g_{k}^{n}\rangle\mathbf{T}_{m,n}^{*},$ (9)
@@ III-E Training Objectives
where $B$ is the batch size of image-text pairs, and $y_{i}$ denotes the ground-truth label of the input image $x_{i}$.
IV Experiments
TABLE II: Comparison with CLIP, CoOp and CoCoOp in the base-to-novel generalization setting. The results demonstrate the strong generalizability to novel classes of our MAP. HM: Harmonic mean to highlight the generalization trade-off [63]. The best results in each column are shown in bold font.
(a) Average results
Method Base Novel HM
CLIP 69.34 74.22 71.70
CoOp 82.69 63.22 71.66
CoCoOp 80.47 71.69 75.83
Ours 83.66 75.76 79.36
(d) DTD
Method Base Novel HM
CLIP 53.24 59.90 56.37
CoOp 79.44 41.18 54.24
CoCoOp 77.01 56.00 64.85
Ours 82.63 66.23 73.53
(g) OxfordPets
Method Base Novel HM
CLIP 91.17 97.26 94.12
CoOp 93.67 95.29 94.47
CoCoOp 95.20 97.69 96.43
Ours 95.43 96.90 96.16
(j) Food101
Method Base Novel HM
CLIP 90.10 91.22 90.66
CoOp 88.33 82.26 85.19
CoCoOp 90.70 91.29 90.99
Ours 90.30 89.30 89.80
(b) ImageNet
Method Base Novel HM
CLIP 72.43 68.14 70.22
CoOp 76.47 67.88 71.92
CoCoOp 75.98 70.43 73.10
Ours 76.60 70.60 73.48
(e) EuroSAT
Method Base Novel HM
CLIP 56.48 64.05 60.03
CoOp 92.19 54.74 68.69
CoCoOp 87.49 60.04 71.21
Ours 92.13 76.10 83.33
(h) StanfordCars
Method Base Novel HM
CLIP 63.37 74.89 68.65
CoOp 78.12 60.40 68.13
CoCoOp 70.49 73.59 72.01
Ours 76.70 73.73 75.18
(k) FGVCAircraft
Method Base Novel HM
CLIP 27.19 36.29 31.09
CoOp 40.44 22.30 28.75
CoCoOp 33.41 23.71 27.74
Ours 41.63 36.43 38.84
(c) Caltech101
Method Base Novel HM
CLIP 96.84 94.00 95.40
CoOp 98.00 89.81 93.73
CoCoOp 97.96 93.81 95.84
Ours 98.30 93.80 96.00
(f) UCF101
Method Base Novel HM
CLIP 70.53 77.50 73.85
CoOp 84.69 56.05 67.46
CoCoOp 82.33 73.45 77.64
Ours 86.67 78.77 82.52
(i) Flowers102
Method Base Novel HM
CLIP 72.08 77.80 74.83
CoOp 97.60 59.67 74.06
CoCoOp 94.87 71.75 81.71
Ours 97.57 75.23 84.95
(l) SUN397
Method Base Novel HM
CLIP 69.36 75.35 72.23
CoOp 80.60 65.89 72.51
CoCoOp 79.74 76.86 78.27
Ours 82.33 76.30 79.20
In this section, we begin by introducing the benchmark settings and implementation details, followed by a comprehensive presentation of the experimental results.
All the models used are based on the open-source CLIP [1] model. We evaluate the adaptation and generalization capability of MAP in four distinct settings following previous methdos [18, 16].
Base-to-novel generalization. Datasets are split into base and novel classes. The model is trained on the training dataset, which is constructed by randomly selecting 16 images per class from base classes. Then the model is evaluated on both base and novel classes. The evaluation encompasses 11 image recognition datasets, including Food101 (Foo) [64], DTD [65], ImageNet (Img) [66], Caltech101 (Cal) [67], EuroSAT (Eur) [68], StanfordCars (Car) [69], FGVCAircraft (FGV) [70], Flowers102 (Flo) [71], OxfordPets (Pet) [72], UCF101 (UCF) [72], and SUN397 (SUN) [73].
Few-shot image classification. To evaluate the learning capacity under extremely limited supervision, we assess the model’s performance across varying shot scenarios, namely, 1, 2, 4, 8, and 16 shots. Similar to the base-to-novel generalization setting, we employ the same 11 datasets.
Domain generalization. To assess the robustness under domain shifts, we train the model using the source dataset ImageNet and subsequently evaluate its performance on out-of-distribution target datasets, namely ImageNet-R (-R) [74], ImageNet-A (-A) [75], ImageNetV2 (V2) [76], and ImageNet-Sketch (-S) [77].
Cross-dataset evaluation. In the cross-dataset transfer setting, we train the models on the source dataset ImageNet and directly evaluate them on target datasets. Specifically, the target datasets include Food101, DTD, Caltech101, EuroSAT, StanfordCars, FGVCAircraft, Flowers102, OxfordPets, UCF101, and SUN397.